Jamal deposited $600 into a savings account that will earn 2 percent simple interest.If Jamal keeps the money in the account for 2 years without making any deposits or withdrawals,how much money will he earn in interest.the second question is how much money will he have in his account including interest after 2 years?
Answer by stanbon(75887)

Jamal deposited $600 into a savings account that will earn 2 percent simple interest.If Jamal keeps the money in the account for 2 years without making any deposits or withdrawals,how much money will he earn in interest.
I = Prt
Interest = 600*0.02*2 = $24.00
--------------------------------------------
the second question is how much money will he have in his account including interest after 2 years?
A = P + Prt
A(2) = 600 + 24 = $624
=================================
Cheers,
Stan H.
---------------------------